The mass and radius of a neutron star is 2MΘ and 15 km, respectively. Calculate the
value of gravitational red shift for light of wavelength 6000 Å at a distance of 2 m
from its surface.
1
Expert's answer
2019-04-26T11:16:54-0400

The gravitational red shift 

"\Delta \lambda=\lambda_0\left(\frac{1}{\sqrt{1-\frac{2GM}{Rc^2}}}-1\right)"

"\Delta \lambda=6000\left(\frac{1}{\sqrt{1-\frac{2\times 6.67\times 10^{-11}\times 4\times10^{30}}{15000\times(3\times 10^8)^2}}}-1\right)=1716 \ A"